NATURAL TRANSITION ORBITALS

RICHARD L. MARTIN, Theoretical Division, Los Almaos National Laboratory, Los Alamos, NM 87545.

A means of finding a compact orbital representation for the electronic transition density matrix is described. The technique utilizes the corresponding orbital transformation of Amos and Hall and allows a dramatic simplification in the qualitative description of an electronic transition.